Your 210 lb (95 kg) patient can tolerate on a plasma concentration of 0.15 mg/L of hydrocoodone only. The Volume of distribution for hydrocofone  is 6 L/Kg. Calculation the dose required for this patient.
Answer #1

Solution: Given, concentration of drug in plasma = Cp = 0.15 mg/L

VD = 6L/kg * 95 kg = 570 L

Now, VD = DB/CP where DB is the amount of drug required.

So, DB = 0.15*570 = 85.5 mg which is the dose required for the patient.